# gamecenter,boxed

### v1

@@ Abstract.
We study the problem of privacy-preserving $k$-means clustering in the horizontally federated setting. Existing federated approaches using secure computation, suffer from substantial overheads and do not offer output privacy. At the same time, differentially private (DP) $k$-means algorithms assume a trusted central curator and do not extend to federated settings. Naively combining the secure and DP solutions results in a protocol with impractical overhead. Instead, our work provides enhancements to both the DP and secure computation components, resulting in a design that is faster, more private, and more accurate than previous work. By utilizing the computational DP model, we design a lightweight, secure aggregation-based approach that achieves four orders of magnitude speed-up over state-of-the-art related work. Furthermore, we not only maintain the utility of the state-of-the-art in the central model of DP, but we improve the utility further by taking advantage of constrained clustering techniques.
Differential Privacy, Federated Learning, $k$-Means Clustering, Secure Computation
ccs: Computing methodologies Cluster analysisccs: Security and privacy Privacy-preserving protocolsccs: Security and privacy Data anonymization and sanitization
1. Introduction
@@
Unsupervised learning allows data analysts to extract meaningful patterns from raw data that may be difficult to label. The canonical example of unsupervised learning is the Euclidean $k$-means clustering problem, where data is grouped into clusters with similar features. Clustering has a plethora of important applications in recommendation systems, fraud detection, and healthcare analytics. In such applications, the dataset often contains sensitive data, which necessitates the use of privacy-preserving techniques. Combining data from multiple parties in the federated scenario yields more robust insights about the global population at the cost of further exasperating the privacy risks.
@@
To counteract the privacy risks, a number of works focus on solving the federated $k$-means problem using secure computation (Vaidya and Clifton, 2003; Jha et al., 2005; Jagannathan and Wright, 2005; Jagannathan et al., 2010; Gheid and Challal, 2016; Bunn and Ostrovsky, 2007; Jäschke and Armknecht, 2018; Mohassel et al., 2020; Patel et al., 2012; Yuan and Tian, 2019; Rao et al., 2015; Xiaoyan et al., 2015; Silva and Bellala, 2017; Jiang et al., 2020). Secure computation enables analysts to solve the $k$-means problem while keeping the sensitive data encrypted. However, all of these works still output the exact result of the clustering. This gives a false sense of privacy; although the input and intermediate computations all remain private, no effort is made to protect the privacy of the output. It is well known in the literature that publishing exact statistics can leak significant information about the input data and allow attacks such as dataset reconstruction (Dwork et al., 2017). This level of privacy does not justify the significant runtime overhead incurred by solely using secure computation.
A separate line of research considers using differential privacy (DP) to protect the privacy of the output of $k$-means (Blum et al., 2005; Dwork, 2011; Su et al., 2016; Mohan et al., 2012; Zhang et al., 2013; Balcan et al., 2017; Lu and Shen, [n. d.]; Park et al., 2017; Ni et al., 2021). Differential privacy guarantees that the clustering’s output will be approximately the same regardless of any individual user’s participation. However, existing methods do not consider the federated setting and assume the existence of a trusted central party who runs the $k$-means algorithm and sanitizes the results with noise before publishing.
@@
To summarize, no prior work offers a solution to protect both the input and output privacy of the federated $k$-means problem. A straw man solution to our problem would be to combine the current state-of-the-art solutions in secure federated $k$-means and DP $k$-means. However, this solution would be unacceptable as the runtime of current federated $k$-means approaches using secure computation is prohibitively large. The state-of-the-art approach of Mohassel et al. (Mohassel et al., 2020) incurs runtime overheads on the order of tens of minutes. Naively adding differentially private noise to this computation will further increase this runtime.
@@
Our work focuses on designing an efficient, DP $k$-means in the federated setting that protects the privacy of the input, output, and intermediate computations. Following related work in federated $k$-means (Mohassel et al., 2020; Jiang et al., 2020), we use Lloyd’s algorithm as the foundation of our protocol. However, we design a new variant of DP-Lloyd (Blum et al., 2005; Su et al., 2016) that noises the centroids directly. Our algorithm improves the clustering utility over the state-of-the-art DP algorithm of Su et al. (Su et al., 2016) by adding constraints to the $k$-means objective (Bennett et al., 2000). Specifically, by adding minimum and maximum constraints to the cluster size, we achieve a non-trivial bound on the sensitivity of the centroids. This enables our algorithm to aggregate at the centroid level and apply less DP noise than prior work that aggregates the sums and counts separately (Su et al., 2016).
@@
To tailor our DP-Lloyd algorithm to secure computation, we design a protocol that publishes the centroids in each step. This allows operations that would be expensive in secure computation (such as assignment and division) to be conducted in plain text. However, unlike prior work (Gheid and Challal, 2016; Jiang et al., 2020), we ensure that the intermediate centroids are protected by DP guarantees. We prove that the protocol is secure in the computational DP definition (Mironov et al., 2009), which allows DP-bounded leakage during the computation. Typically, utilizing the computational DP model to leak intermediate steps implies sacrificing privacy or utility for an efficient runtime. However, in our case, we specifically choose to leak intermediate values that are typically accounted for in the central DP proof of DP-Lloyd (Blum et al., 2005; Su et al., 2016). In other words, we get this speed-up with no additional cost to the privacy parameter $\epsilon$. Furthermore, rather than decrease the utility, we improve the utility of the clustering over state-of-the-art approaches (Su et al., 2016).
@@
To perform the aggregation over multiple parties, we design a modified secure aggregation protocol that keeps the aggregator oblivious to the client’s inputs and the centroids that are output at each iteration of the protocol. This allows us to add DP noise equivalent to the central model of DP. The combination of our improved DP algorithm, leaking the DP centroids, and our lightweight secure aggregation protocol allows us to reduce the computation time by up to four orders of magnitude compared to related work (Mohassel et al., 2020) (from minutes to a fraction of a second per iteration).
In summary, our contributions include:
@@
  • We improve the clustering utility over state-of-the-art DP $k$-means solutions by using cluster-size constraints and centroid-level aggregation.
  • We design an efficient protocol using DP and a modified secure aggregation protocol to implement our protocol in the local model of trust.
  • We prove our protocol is secure and preserves the end-to-end privacy in the computational model of DP and reduces the runtime by four orders of magnitude over the state-of-the-art secure federated approaches.
The remainder of the paper is organized as follows. We begin with some problem-specific background information in Section 2 and formally define the problem in Section 3. In Section 4, we summarize the relevant literature. We then present our complete protocol for federated DP $k$-means (FastLloyd) and prove its privacy and utility in Section 5. In Section 6, we give an in-depth evaluation of our protocol in terms of utility, runtime, and communication size over various real-world and synthetic datasets. Finally, Section 7 describes extensions to our protocol, including a distance-constrained method and alternatives to the aggregation server.
2. Background
2.1. Notation
Throughout this paper, we will denote objects that we intend to further slice/index with boldface notation (e.g. $\boldsymbol{\mu}$, while keeping atomic objects (whether scalars or vectors) as normal face (e.g. $\epsilon$). A summary of the notation used in this paper is provided in Table 1 and we will define the notation as it is used.
@@ 2.1. Notation
$\boldsymbol{\mu}$ Centroids of clusters: $\{\mu_{1},\mu_{2},...,\mu_{k}\}$
$\boldsymbol{\gamma}$ Differentially-private noise
$\eta$ Constraint parameter
$\epsilon$ Privacy budget
Table 1. Notation used in the paper.
2.2. $k$-means Lloyd’s Algorithm
The $k$-means problem is a discrete optimization problem that aims to partition a set of $N$ $d$-dimensional observations into $k$ clusters, each represented by its mean or centroid. Given a dataset of observations $D=\{x_{1},x_{2},...,x_{N}\}$, where each observation is a $d$-dimensional real vector, the $k$-means problem is to find an assignment of data points to clusters, and a set of cluster centers, that minimizes the Within-cluster Sum of Squares (WCSS) objective:
(1) $\underset{\mathbf{O},\boldsymbol{\mu}}{\text{argmin}}\sum_{j=1}^{k}\sum_{x_{l}\in O_{j}}||x_{l}-\mu_{j}||^{2}$
where $\mathbf{O}=\{O_{1},O_{2},...,O_{k}\}$, $O_{\subseteq}D$ are the clusters, and $\boldsymbol{\mu}=\{\mu_{1},\mu_{2},...,\mu_{k}\}$ are the centroids of the clusters, defined to be the (arithmetic) mean of points in $O_{j}$.
Lloyd’s Algorithm
The $k$-means problem is NP-hard in Euclidean space (Aloise et al., 2009). However, a simple heuristic known as Lloyd’s algorithm is commonly applied for practical applications (Lloyd, 1982). We detail the basic procedure of Lloyd’s algorithm:
  1. (1)
@@ 2.2. $k$-means Lloyd’s Algorithm
      $\mu_{j}^{(t)}=\frac{\sum\limits_{x_{l}\in\mathbf{O}{j}^{(t)}}x_{l}}{\left|O_{j}^{(t)}\right|}$
2.3. Constrained $k$-means
The constrained $k$-means problem is a variant of the $k$-means problem where, in addition to minimizing the WCSS (Eqn 1), a certain set of constraints $\boldsymbol{\eta}$ must be satisfied. The problem was first introduced by Bennet et. al (Bennett et al., 2000), where they propose a method to enforce minimum cluster size constraints, $\boldsymbol{\eta}^{min}$.
To solve the constrained $k$-means problem, Lloyd’s algorithm can be modified to incorporate the constraints in the assignment step of every iteration. However, the assignment step becomes nontrivial as it requires the solution to a constrained optimization problem. Bennet et. al (Bennett et al., 2000) provide an efficient solution to this constrained optimization problem by formulating it as a min-cost flow problem, which can be solved using off-the-shelf network flow solvers. An open-source implementation by Levy-Kramer (Levy-Kramer, 2018) implements the approach, using Google’s ORTools (Perron and Furnon, 2023) to solve the flow problem, and extends it to allow for maximum cluster size constraints, $\boldsymbol{\eta}^{max}$ (Eqn 2).
(2) $\underset{\mathbf{O},\boldsymbol{\mu}}{\text{argmin}}\sum_{j=1}^{k}\sum_{x_{l}\in O_{j}}||x_{l}-\mu_{j}||^{2}\quad\text{s.t.}\quad\eta_{j}^{min}\leq|O_{j}|\leq\eta_{j}^{max}\quad\forall j$
Figure 1 shows an example reduction of an assignment instance to a network flow graph for $\boldsymbol{\eta}^{min}$ and $\boldsymbol{\eta}^{max}$.
[…]
Figure 1. An example reduction of an assignment instance ($N=5,k=3$) to a network flow graph for $\boldsymbol{\eta}^{min}$ and $\boldsymbol{\eta}^{max}$.
Datapoints are represented as nodes and connected to cluster nodes through directed edges. The cost of each edge $(x_{l}\rightarrow O_{j})$ is defined to be the Euclidean distance between $x_{l}$ and the corresponding centroid $\mu_{j}$. The capacity of each edge is set to $1$, and each data node has a supply of $1$, ensuring that each data point is associated with exactly one cluster. On the other hand, each centroid node has a supply of $-\eta_{min}$ (equivalently, a demand of $\eta_{min}$). As every valid flow solution must satisfy this demand, a minimum of $\eta_{min}$ data points will flow into each cluster. To impose the maximum size constraint, clone nodes are created for each centroid, and each centroid is connected with its clone through an edge with zero cost and a capacity of $\eta_{max}$. This constrains the maximum flow into each cluster to be $\eta_{max}$. Lastly, a sink node is introduced to consume the remaining flow. A valid flow solution respects the constraints on cluster sizes, and by minimizing the cost, the total distance of points to their assigned centroids is minimized. Therefore, a min-cost flow solution to this model solves the constrained $k$-means problem for both minimum and maximum cluster constraints.
2.4. Differential Privacy
Differential privacy (DP) (Dwork et al., 2006) is an increasingly popular notion for protecting individuals’ privacy while allowing the computation of aggregate statistics. Differential privacy guarantees that an algorithm’s output is approximately the same, regardless of the participation of any single user. More formally, differential privacy can be defined as follows.
@@ 2.4. Differential Privacy
A randomized algorithm $M:\mathcal{D}\mapsto\mathbb{R}$ is $\epsilon$-DP, if for any pair of neighbouring datasets $D,D^{\prime}\in\mathcal{D}$, and for any $T\subseteq\mathbb{R}$ we have
(3) $\Pr[M(D)\in T]\leq e^{\epsilon}\Pr[M(D^{\prime})\in T].$
The privacy parameter $\epsilon$ defines how similar the outputs must be and determines the trade-off between privacy and utility. We use the unbounded neighbouring definition where datasets are neighbours if $|D\backslash D^{\prime}\cup D^{\prime}\backslash D|=1$. That is, we allow the addition or removal of a single data point. We note that arbitrary computations can be carried out on the output of a DP mechanism without affecting privacy (the postprocessing lemma (Dwork and Roth, 2014)). Finally, DP is composed naturally with multiple runs of a mechanism. If we apply a differentially private mechanism(s) sequentially, the privacy parameters are composed through summation or more advanced methods (Dwork and Roth, 2014). If a mechanism is applied multiple times over disjoint subsets of the dataset, then the total privacy leakage is the maximum privacy parameter over each subset (parallel composition (Dwork and Roth, 2014)).
@@
(4) $\Delta^{(f)}_{d}=\max_{(D,D^{\prime})}\mathbb{D}(f(D),f(D^{\prime})),$
@@
Common examples of the distance metric are the $\ell_{1}$-norm and $\ell_{2}$-norm. With the notion of sensitivity, we can now define the Laplace mechanism, which uses sensitivity based on the $\ell_{1}$-norm, which we denote $\Delta^{(f)}_{1}$ for simplicity.
Definition 2.3 (Laplace Mechanism).
Let $f:\mathcal{D}\mapsto\mathbb{R}^{k}$. The Laplace mechanism is defined as
(5) $M(D)=f(D)+\boldsymbol{\gamma},$
where
$\boldsymbol{\gamma}\in\mathbb{R}^{k}\quad\text{s.t.}\quad\gamma_{i}\overset{\mathrm{iid}}{\sim}Lap(\Delta^{(f)}_{1}/\epsilon)$
i.e. $\boldsymbol{\gamma}$ represents $k$ i.i.d draws from the Laplace distribution with scale parameter $\Delta^{(f)}_{1}/\epsilon$.
It is well known that the Laplace mechanism satisfies $\epsilon$-DP (Dwork and Roth, 2014).
3. Problem Statement
We consider $M$ clients denoted by $\boldsymbol{p}=\{p_{1},p_{2},...,p_{M}\}$, with each party $p_{i}$ owning a private dataset $D_{i}=\{x_{1},x_{2},...,x_{N_{i}}\}$, where $x_{l}\in\mathbb{R}^{d}$, $d$ denotes the dimensionality of the dataset and $N_{i}$ is the size of the dataset of party $i$. For simplicity, we assume that datasets held by each client are of similar sizes. The objective is to compute a collaborative $k$-means clustering $(\mathbf{O},\boldsymbol{\mu})$ over $\boldsymbol{D}$ that minimizes the WCSS objective (1); where $\boldsymbol{D}=\bigcup_{i=1}^{i=M}D_{i}$ is the union of the datasets held by the clients, $\mathbf{O}=\{O_{1},O_{2},...,O_{k}\}$ are the clusters, and $\boldsymbol{\mu}=\{\mu_{1},\mu_{2},...,\mu_{k}\}$ are the respective centroids of the clusters.
We aim to optimize this objective while formally proving the privacy of the output and intermediate computations. Specifically, the protocol should be secure in the computational model of differential privacy introduced by Mirnov et al. (Mironov et al., 2009) and extended to the multi-party setting by Humphries et al. (Humphries et al., 2022).
Definition 3.1 (IND-CDP-MPC (Humphries et al., 2022; Mironov et al., 2009)).
A multi-party protocol $\Pi$ for computing function $f$ satisfies $\epsilon(\lambda)$- indistinguishable computationally differential privacy (IND-CDP-MPC) if for every probabilistic polynomial time (in $\lambda$) adversary $\mathcal{A}$ with input $D_{A}$, and for neighbouring datasets $D,D^{\prime}$ belonging to the honest parties (i.e. $D,D^{\prime}=\cup_{i\setminus A}D_{i}$),

### v2

@@ Abstract
We study the problem of privacy-preserving $k$-means clustering in the horizontally federated setting. Existing federated approaches using secure computation suffer from substantial overheads and do not offer output privacy. At the same time, differentially private (DP) $k$-means algorithms either assume a trusted central curator or significantly degrade utility by adding noise in the local DP model. Naively combining the secure and central DP solutions results in a protocol with impractical overhead. Instead, our work provides enhancements to both the DP and secure computation components, resulting in a design that is faster, more private, and more accurate than previous work. By utilizing the computational DP model, we design a lightweight, secure aggregation-based approach that achieves five orders of magnitude speed-up over state-of-the-art related work. Furthermore, we not only maintain the utility of the state-of-the-art in the central model of DP, but we improve the utility further by designing a new DP clustering mechanism.
1 Introduction
@@
Unsupervised learning allows data analysts to extract meaningful patterns from raw data that may be difficult to label. The canonical example of unsupervised learning is the Euclidean $k$-means clustering problem, where data is grouped into clusters with similar features. Clustering has a plethora of important applications in recommendation systems, fraud detection, and healthcare analytics [30]. In such applications, the dataset often contains sensitive data, which necessitates the use of privacy-preserving techniques. Combining databases horizontally split across multiple parties in the federated scenario yields more robust insights about the global population at the cost of further exasperating the privacy risks.
@@
To counteract the privacy risks, a number of works focus on solving the federated $k$-means problem using secure multi-party computation [66, 37, 35, 34, 29, 11, 36, 50, 56, 70, 59, 69, 60, 38]. Secure computation enables analysts to solve the $k$-means problem while keeping the sensitive data encrypted. However, all of these works still output the exact result of the clustering. This gives a false sense of privacy; although the input and intermediate computations all remain private, no effort is made to protect the privacy of the output. It is well known in the literature that publishing exact statistics can leak significant information about the input data and allow attacks such as dataset reconstruction [19]. This level of privacy does not justify the significant runtime overhead incurred by solely using secure computation.
A separate line of research considers using differential privacy (DP) to protect the data. Differential privacy guarantees that the output (the cluster centroids) will be approximately the same regardless of any individual user’s participation. Differential privacy can be applied in the central, local, or shuffle models. In the central model, a trusted aggregator receives the input as plaintext (unprotected) and randomizes the cluster centroids to ensure the output satisfies DP. Several works [9, 16, 64, 49, 71, 3, 45, 55, 52] have considered the central setting. In the local model, instead of using a trusted aggregator, each party perturbs their data locally before sending it to the aggregator. While this removes the need for a trusted aggregator (similar to secure computation), the utility at high privacy levels is low. Specifically, the utility is asymptotically worse than the central model by a factor of $\sqrt{n}$, where $n$ is the number of data points [12]. Several works [54, 62, 13] have considered the local model. The shuffle model is a hybrid between the local and central models, with two mutually distrustful parties: a shuffler and an aggregator. Despite improving over the local model, the shuffle model still incurs a $6]{n}$ degradation in utility [5].
@@
To summarize, no prior work offers a solution that protects both the input and output privacy of the federated $k$-means problem and provides a good utility vs. privacy trade-off. A straw man solution to our problem would be to combine the current state-of-the-art solutions in federated $k$-means using secure computation and $k$-means in the central DP model. However, this solution would be unacceptable as the runtime of current federated $k$-means approaches using secure computation is prohibitively large. The state-of-the-art approach of Mohassel et al. [50] incurs runtime overheads on the order of tens of minutes. Naively adding differentially private noise to this computation will further increase this runtime.
@@
Our work focuses on designing an efficient DP $k$-means in the federated setting (using secure computation) that protects the privacy of the input, output, and intermediate computations. Following related work in federated $k$-means [50, 38], we use Lloyd’s algorithm as the foundation of our protocol. However, we design a new variant of DP-Lloyd [9, 64] that achieves a tighter bound on the sensitivity. Our algorithm significantly improves the clustering utility over the state-of-the-art DP algorithm of Su et al. [64], especially in higher dimensions and number of clusters.
Our algorithm is of independent interest in the central model of DP. First, we enforce a bound on the radius of each cluster in the assignment step of Lloyd’s algorithm. We then modify Lloyd’s algorithm to compute updates relative to the previous centroid. Our relative updates have a sensitivity proportional to the bounded cluster radius rather than the domain size used in previous work [9, 64]. We find that our radius-based sensitivity bound is best suited to the use of the analytic Gaussian mechanism [6, 15], which further improves utility. Finally, we utilize additional post-processing steps based on the radius and the domain to improve the algorithm further. Together, these components lead to an improvement of up to 88% in utility (reduction in clustering error).
@@
To tailor our DP-Lloyd algorithm to secure computation, we design a protocol that publishes intermediate computations in each iteration. This allows operations that would be expensive in secure computation (such as assignment and division) to be conducted in plain text. However, unlike prior work [29, 38], we ensure that the intermediate computations are protected by DP guarantees. We prove that the protocol is secure in the computational DP definition [48], which allows DP-bounded leakage during the computation. Typically, utilizing the computational DP model to leak intermediate steps implies sacrificing privacy or utility for an efficient runtime. However, in our case, we specifically choose to leak intermediate values already accounted for in the central DP proof of DP-Lloyd [9, 64]. In other words, we get this speed-up with no additional cost to the privacy parameter $\epsilon$. Furthermore, rather than decrease the utility, we improve the utility of the clustering over state-of-the-art approaches [64].
@@
To perform the aggregation over multiple parties, we design a lightweight secure aggregation protocol that keeps the aggregator oblivious to the client’s inputs and the global centroids that are output at each iteration of the protocol. This allows the aggregator to add DP noise equivalent to the central model of DP. The combination of our improved DP algorithm, leaking the DP centroids, and our lightweight, secure aggregation protocol allows us to reduce the computation time by up to five orders of magnitude compared to related work [50] (from minutes to milliseconds per iteration).
In summary, our contributions are four-fold:
@@
  • We improve the clustering utility over state-of-the-art DP $k$-means solutions by developing a new DP algorithm with various improvements, such as enforcing a radius constraint on the centroids and using relative cluster updates.
  • We design an efficient protocol using DP and a lightweight secure aggregation protocol to implement our protocol in the local trust model.
  • We prove our protocol is secure and preserves the end-to-end privacy in the computational model of DP and reduces the runtime by five orders of magnitude over the state-of-the-art secure federated approaches.
The remainder of the paper is organized as follows. We begin with some problem-specific background information in Section 2 and formally define the problem in Section 3. In Section 4, we summarize the relevant literature. We then present our complete protocol for federated DP $k$-means (FastLloyd) and prove its privacy and utility in Section 5. Finally, in Section 6, we give an in-depth evaluation of our protocol in terms of utility, runtime, and communication size over various real-world and synthetic datasets.
2 Background
2.1 Notation
Throughout this paper, we will denote objects that we intend to further slice/index with boldface notation (e.g. $\bm{\mu}$), while keeping atomic objects (whether scalars or vectors) as normal face (e.g. $\epsilon$). A summary of the notation used in this paper is provided in Table 1, and we will define the notation as it is used.
@@ 2.1 Notation
$\bm{\mu}$ Centroids of clusters: $\{\mu_{1},\mu_{2},...,\mu_{k}\}$
$S_{j}$ Sum of data points in cluster $j$
$C_{j}$ Count of data points in cluster $j$
$R_{j}$ Relative sum of data points in cluster $j$
$\bm{\gamma}$ Differentially-private noise
$\eta$ Maximum radius bound
$(\epsilon,\delta)$ Privacy budget
$\sigma$ Noise multiplier
Table 1: Notation used in the paper.
2.2 $k$-Means Problem
The $k$-means problem is a discrete optimization problem that aims to partition a set of $N$ $d$-dimensional observations into $k$ clusters, each represented by its mean or centroid. Given a dataset of observations $D=\{x_{1},x_{2},...,x_{N}\}$, where each observation is a $d$-dimensional real vector, the $k$-means problem is to find an assignment of data points to clusters, and a set of cluster centroids, that minimizes the Within-cluster Sum of Squares (WCSS) objective:
$\underset{\mathbf{O},\bm{\mu}}{\text{argmin}}\sum_{j=1}^{k}\sum_{x_{l}\in O_{j}}||x_{l}-\mu_{j}||^{2}_{2}$ (1)
where $\mathbf{O}=\{O_{1},O_{2},...,O_{k}\}$, $O_{j}\subseteq D$ are the clusters, and $\bm{\mu}=\{\mu_{1},\mu_{2},...,\mu_{k}\}$ are the centroids of the clusters, defined to be the (arithmetic) mean of points in $O_{j}$.
2.2.1 Lloyd’s Algorithm
The $k$-means problem is NP-hard in Euclidean space [2]. However, a simple heuristic known as Lloyd’s algorithm is commonly applied for practical applications [44]. We detail the basic procedure of Lloyd’s algorithm:
  1. 1.
@@ 2.2.1 Lloyd’s Algorithm
      $\mu_{j}^{(t)}=\frac{\sum\limits_{x_{l}\in O_{j}^{(t)}}x_{l}}{\left|O_{j}^{(t)}\right|}$
2.3 Differential Privacy
Differential privacy (DP) [17] is an increasingly popular notion to protect the privacy of individuals while allowing the computation of aggregate statistics. Differential privacy guarantees that an algorithm’s output is approximately the same, regardless of the participation of any single user. More formally, differential privacy can be defined as follows.
@@ 2.3 Differential Privacy
A randomized algorithm $M:\mathcal{D}\mapsto\mathbb{R}$ is $(\epsilon,\delta)$-DP, if for any pair of neighbouring datasets $D,D^{\prime}\in\mathcal{D}$, and for any $T\subseteq\mathbb{R}$ we have
$\Pr[M(D)\in T]\leq e^{\epsilon}\Pr[M(D^{\prime})\in T]+\delta.$ (2)
The privacy parameter $\epsilon$ defines how similar the outputs must be, and $\delta$ allows a small chance of failure in the definition. We use the unbounded neighbouring definition where datasets are neighbours if $|D\backslash D^{\prime}\cup D^{\prime}\backslash D|=1$. That is, we allow for the addition or removal of a single data point. We note that arbitrary computations can be carried out on the output of a DP mechanism without affecting privacy (the post-processing lemma [18]). Finally, DP is composed naturally with multiple runs of a mechanism. If we apply a differentially private mechanism(s) sequentially, the privacy parameters are composed through summation or more advanced methods [18]. If a mechanism is applied multiple times over disjoint subsets of the dataset, then the total privacy leakage is the maximum privacy parameter over each subset (parallel composition [18]).
@@
$\Delta^{(f)}=\max_{(D,D^{\prime})}\mathbb{D}(f(D),f(D^{\prime})),$ (3)
@@
We will focus on the $\ell_{2}$ norm in this work as we use the Gaussian Mechanism. To analyze the Gaussian Mechanism, we will use Gaussian Differential Privacy (GDP) [15].
Definition 2.3 (GDP [15]).
A mechanism $M$ is said to satisfy $\theta$-Gaussian Differential Privacy ($\theta$-GDP) if it is $G_{\theta}$-DP. That is,
$T\big{(}M(D),M(D^{\prime})\big{)}\geq G_{\theta}$
for all neighboring datasets $D$ and $D^{\prime}$, where $T$ is a trade-off function and $G_{\theta}=T\big{(}\mathcal{N}(0,1),\mathcal{N}(\theta,1)\big{)}$ (see Dong et al. [15] for more details).
Naturally, the Gaussian Mechanism satisfies GDP.
Theorem 2.1.
(Gaussian Mechanism GDP [15]) Define the Gaussian mechanism that operates on a statistic $f$ as $M(D)=f(D)+\gamma$, where $\gamma\sim\mathcal{N}(0,(\Delta^{(f)})^{2}/\theta^{2})$. Then, $M$ is $\theta$-GDP.
Similar to DP, GDP composes over multiple adaptive uses of a mechanism.
Theorem 2.2 (GDP Composition [15]).
The $n$-fold composition of $\theta_{i}$-GDP mechanisms is $\sqrt{\theta_{1}^{2}+\cdots+\theta_{n}^{2}}$-GDP.
Finally, it is possible to convert a GDP guarantee to DP and vice versa:
Theorem 2.3 (GDP to DP [15, 6]).
A mechanism is $\theta$-GDP if and only if it is $\big{(}\epsilon,\delta(\epsilon)\big{)}$-DP for all $\epsilon\geq 0$, where
$\delta(\epsilon)=\Phi\Big{(}-\frac{\varepsilon}{\theta}+\frac{\theta}{2}\Big{)}-e^{\epsilon}\Phi\Big{(}-\frac{\varepsilon}{\theta}-\frac{\theta}{2}\Big{)}.$
In practice, we use the algorithm derived by Balle and Wang to solve this function for $\theta$ [6, Algorithm 1].
3 Problem Statement
We consider $M$ clients denoted by $\bm{p}=\{p_{1},p_{2},...,p_{M}\}$, with each party $p_{i}$ owning a private dataset $D_{i}=\{x_{1},x_{2},...,x_{N_{i}}\}$, where $x_{l}\in\mathbb{R}^{d}$, $d$ denotes the dimensionality of the dataset and $N_{i}$ is the size of the dataset of party $i$. The total dataset that we compute on, can be very large, even if there are only a small number of clients. The objective is to compute a collaborative $k$-means clustering $(\mathbf{O},\bm{\mu})$ over $\bm{D}$ that minimizes the WCSS objective (1); where $\bm{D}=\bigcup_{i=1}^{i=M}D_{i}$ is the union of the datasets held by the clients, $\mathbf{O}=\{O_{1},O_{2},...,O_{k}\}$ are the clusters, and $\bm{\mu}=\{\mu_{1},\mu_{2},...,\mu_{k}\}$ are the respective centroids of the clusters.
We aim to optimize this objective while formally proving the privacy of the output and intermediate computations. Specifically, the protocol should be secure in the computational model of differential privacy introduced by Mirnov et al. [48] and extended to the multi-party setting by Humphries et al. [33]. We further extend this model to the approximate DP setting by incorporating the failure probability $\delta$.
Definition 3.1 (IND-CDP-MPC [33, 48]).
A multi-party protocol $\Pi$ for computing function $f$ satisfies $(\epsilon(\lambda),\delta)$- indistinguishable computationally differential privacy (IND-CDP-MPC) if for every probabilistic polynomial time (in $\lambda$) adversary $\mathcal{A}$ with input $D_{A}$, and for neighbouring datasets $D,D^{\prime}$ belonging to the honest parties (i.e. $D,D^{\prime}=\cup_{i\setminus A}D_{i}$),